Research Quick look 95W Data display Data delivery TAO Home
TAO/EPIC  logo Status of sensors
Summary of current deployments - History of EPIC sites - TAO/EPIC Cruises along 95W

06 Nov 2003

Dates in the sensor fields are the failure dates for the sensor. If "Tube" field contains a failure date, there will not be realtime data for any sensor at this site after this date. Blank cell indicates sensors are reporting good data. "X" indicates no sensor at that depth, "+" indicates sensor that is sending data intermittantly. Red columns indicate buoy is not transmitting. Yearday calendar

12 N 10 N 8 N 5 N 3.5 N 2 N 0 2 S 5 S 8 S
Depl or repair date 2003/112 2003/111 2003/111 2003/109 2002/289 2003/108 2003/107 2003/106 2003/105 2003/104



















RH 2003/275









T 5m  


C 5m  


T 10m 2002/327 





C 10m 2002/327 
2003/111 x



Velocity 10m 2002/327   x  

T 20m    


C 20m    


T 40m 2003/236  


C 40m     x

Velocity 40m X X X x x x x x x
T 60m    

T 80m     2003/231


C 80m    


T 100m    


T 120m     2003/231


C 120m     x


T 140m     2003/231


T 180m     2003/231

TP 300m     2003/231


TP 500m    



Date = Year/julian day
ATRH = Air temperature / Relative Humidity
SST = Sea Surface Temperature
SSC = Sea Surface Conductivity
SWR = Short wave radiation
LWR = Long wave radiation
BP = Barometric Pressure
T'n' = Temperature module and depth of module
C'n' = Conductivity module and depth of module
TP'n' = Temperature/pressure module and depth of module
Columns in red indicate major mooring failure (not transmitting data or position)

History of EPIC sites

12N 95W
PM123 - Deployed 99/336 - 00/111. Mooring was stripped of all surface instrumentation by vandals (00/068). Subsurface modules and toroid were recoverd.
PM141 - Deployed 00/112 - 01/116. SSC bad since depl. Wind vane 0 (00/201). Tube stopped transmitting 00/304 - no warning signs of failure. REC NOTES - Tower (and all surface instrumentation) was missing, and appeared to have been unbolted and taken.
PM141B- 00/317. Winds failed after depl, C80 failed. 01/022 - all subsurface data 0. REC NOTES: Top section broken loose from nilspin socket. 10m and 180m modules missing. UW acoustic rain gauge hanging by one fistgrip and its cage enclosure was damaged. Buoy and tower sensors heavily covered w/guano.
PM211 - Deployed 01/116 -02/091. V10m failed at deployment. Rain FFFF (01/179). 120m cond failed (01/118)
PM211B - Swapped rain gauge (01/254). Old gauge coated w/guano and feathers on intake. Rain gauge failed after 11 days. 20m cond failed (01/318)
PM211C - Swapped rain gauge (01/324). LWR failed (02/002). REC NOTES: Rain gauge had a smooth solid 2-3 cm level pool of guano inside it, above the screen. No way for water to penetrate. Gauge may not be salvagable.
PM279 - Deployed, 02/091 - 02/285. All subsurface 0 02/111. REC NOTES: Rain gauge clogged w/crusty guano and some electical tape. Dead bird in toroid. Nilspin jacket stripped to metal core at 20m where module was.
PM319A - Deployed 02/285 -03/300 TCV 10m failed 02/327. Winds bad entire "A" depl (see swap notes) SWAP NOTES - Anemometer was mounted incorrectly on mast and winds were flagged after recovery.
PM319B - 03/112 - Swapped rain gauge.RH failed, c40 reads low. REC NOTES: Tower ring broken at weld and bent. 10m TCV missing
and cable broken at connector

10N 95W
PM122 - Deployed 99/335 - 00/316. Subsurface modules intermittently 0. Rain failed (000/90-00/112).
PM122B - 00/112 Swapped tube, SSC, t20, t40, and rain 00/112. REC NOTES - Release did not fire. Cut wire at 200m, and all instrumentation below 200m lost. SWR bent, tail and one blade of prop of anemometer missing. 10m TC missing. Small amt of fishing gear on SST leg of bridle.
PM183 - 00/316 -01/115. C80 failed after depl. Subsurface data intermittantly (frequently) 0 after 01/042. REC NOTES: Plastic jacketing on nilspin was worn down to bare wire just above 12m Sontek. 12m Sontek and 20 m module gone. Lots of guano on met sensors.
PM210-Deployed 01/115 - 02/090. SWR and rain flaky (01/139), flagged.
PM210B - Swapped rain gauge and SWR (01/254). Old rain gauge looked ok, SWR had condensation in dome. Installed 6/10 sbe's for Carter Ohlmann (dive ops). Subsurface intermittantly 0. 5m cond failed (01/273)
PM210C - Swapped rain gauge (01/325). Subsurface working ok. TCV failed (01/332) REC NOTES: Found 3m section of jacketing on nilspin stretched thin enough to expose wire cable underneath a few meters below Sontek. Some inductive loss below this point.
PM 278 - Deployed 02/090 -02/285. All subsurf 0 02/106. Wind speed 0, co/va ok (02/161). REC NOTES:Rain gauge, radiometers, anemometer gone. One mast bent on tower ring atrh mount bent up 70 deg. 3 dead birds in toroid. 5m (12386), 10m (13116), 12m sontek (d56), 20m (12387), 40m (12388) all missing.
PM321A - Deployed 02/286-03/301. Winds bad entire "A" depl (see swap notes) SWAP NOTES- Anemometer was mounted incorrectly on mast and winds were flagged after recovery. Lots of guano in rain gauge.
PM321B - 03/111 -Swapped anemometer and rain gauge. C10 flagged because of density inversion w/surrounding sensors (03/111). Winds failed 03/178 speed 0.SWR failed 03/297 REC NOTES: Anemometer missing, LWR had shield completely missing.

8N 95W
PM121 - Deployed 99/334 - 00/315. SSC/SST, rain failed
PM121B - 00/113 - swapped tube, SSC, T5-T40, rain; added LWR and BP. SSC, wind, ATRH failed, and Tube stopped transmitting simultaneously (00/203). REC NOTES - Tower (and all surface instrumentation) was missing.
PM182 - Deployed 00/315 - 01/326.
PM182B -01/115 .Reset tube clock to GMT (was 5 hours off, fixed in processing so that it was transparent to researchers). LWR failed (01/183). Winds anomalous but possibly real - not flagged (01/199).
PM264 - Deployed 01/326 - 02/287. 10m current failed (01/346).
PM264B - 02/089. Rain swap. Comments: Lots of guano on shelves and on top of ATRH. 30 feet of polypro tied on buoy leg and trailing in water. Some guano inside rain gauge and some small feathers on screen. Lots of guano on top of rain mast. REC NOTES: Rain gauge, barometer guano caked. Plenty of barnacles and algae to 20m. Fishing line around 5m TC11 module.
PM323 - Deployed 02/287-03/111. 10m current failed 02/322. RH failed.REC NOTES: Polypro line around tower and bent atrh, 2" line around base of tower down to 12m sontek, assorted fishing gear down to 100m. Only 3 blades left on anemometer propeller. Heavy duty bird spikes on rain gauge all bent
PM366 - Deployed 03/111-03/302. SSC failed 03/114. Rain erratic 03/118, Wind speed and vane 0 (03/224). Most subsurf 0 03/231, AT and RH 0 03/233. REC NOTES: One fishing vessel at buoy as ship approached. 2nd fishing boat was about 8nm away. Anemometer missing, ATRH missing, and mount broken at tower weld. Abou 50-75' of 1"-2" hawser around buoy and govs of long liner gear around top module and at various depths below that.

5N 95W
PM087B - Deployed 99/333 - 00/114. Winds seemed anomalous (and were flagged) but later were "unflagged" as 5 day analysis looked ok. Observer report (00/062) from anonymous fishing boat - reported boat did 'slingshot' on buoy "buoy returned to its position with no apparent damage".
PM142 - Deployed 00/115 -* . Subsurface modules intermittently 0. Buoy adrift (00/159), stopped transmitting (00/192)
*Buoy never recovered.
PM181 - Deployed 00/314 - 01/327. V10 failed (00/365). Winds anomalous from start.SWAP NOTES: Wind sensor prop bearings were shot, with front one completely disintegrated.
PM181B - 01/114. Swapped wind sensor, reset tube, and ATRH read 0, so swapped ATRH. REC NOTES: Longline gear at 100m. TC1, T5, TCV 10 lost. No data 12m Sontek. 42m Sontek dead. Broken bracket for rain gauge
PM265 - Deployed 01/327- 02/288. 40m cond failed (01/327). Winds anomalous w/climatology and surrounding sites (02/032). REPAIR 02/088 - Orientation pin on mast missing and RMY loose on mast. Replaced pin and reseated RMY (no instrument change). 80m cond failed (02/055). 40m currents failed (02/123). All subsurf 0 02/223. REC NOTES T8 (180m) and T9 (300m) missing. Fishing line wrapped around chain attached to release.
Rain gauge clogged w/guano.
PM324 - Deployed 02/288-03/303 C10 density inversion 2003/064, SWR erratic 03/097. SWAP NOTES: SWR and BP cables cut at fat leg. Floats drifting 200m from mooring.
PM324B Swapped SWR and rain 2003/109. Consistent density inversion at 20m, C20 flagged 2003/111. RH about 20% lower than surrounding sites (2003/151). SWR unrealistic values 03/162. Rain erratic 03/220. Wind speed 0 03/258. REC NOTES: Anemometer gone. Small amount of fishing line on nilspin termination and above Sontek heads. Large wads of fishing line below 60m module and 100m module. Buoy had several large cracks in fiberglass above water line.

3.5N 95W
PM120 - Deployed 99/332 - 00/115. Subsurface modules intermittently 0. No modules, including Sontek on wire above 20m at recovery.
PM143 - Deployed 00/116 -01/113. Wind -> 0 (00/142) REC NOTE: - Prop and shaft of wind sensor were missing, it appeared the set screw holding the magnet and bearing in place had fallen out. No signs of vandalism.
PM143B - 00/313. Swapped anemometer. SSC failed (00/328). Subsurface data intermittantly 0 starting 01/039. REC NOTES: Some longline gear in mooring. Top 5 modules to 40m heavily fouled with barnacles.
PM209 - Deployed 01/113 - 01/328. All subsurface 0 (01/307). REC NOTE: Longline gear at 10-20m.
PM266 - Deployed 01/328-02/087. Stopped transmitting 01/346. Resumed transmissions 01/360. All data bad (looks like major tube damage). Tube transmitting from Ecuador.REC NOTES: Recovered subsurface portion of mooring only. Tower completely stripped, including radar reflector. Piece of poly tied to one leg, float attached. Fishing line around nilspin at 80m. All above waterline instr lost (tube, anemometer, atrh, SWR, LWR, BP, rain).
PM277 - Deployed 02/087 -*. Adrift (02/229) and stopped transmitting 02/232. *Buoy lost. No response from release
PM325 - Deployed 02/289- 03/305 . SSC reads low 03/196. C40 shows consistent density inverstion 03/210. SST failed 2003/293. REC NOTES: Buoy looked good, had fishing line at 70m and again at 140m down past 180m, but not a huge amt. Had small amt just below 500m module, at top of first reel of nylon, and at top of 3rd reel of nylon. Chain and swivel above release had small wad with one line down to bottom of release. Pod of 4 pilot whales feeding around buoy during sensor removal.

2N 95W
PM118 -Deployed 99/331 - 00/312. T/P 500 failed at deployment.
PM118B - 00/116 Swapped tube, 0-40m modules, added LWR and BP. SSC failed (00/158), Wind ->0 (00255).Email from tuna taggers taking CTD's, XBT's near buoy reported mooring had recently been fished (00/259). REC NOTES - release did not work, and wire was cut at ~190m (all instruments below 190m were lost).
PM-180 - Deployed 00/313 - *. SWR and LWR bad since depl. Tube stopped transmitting (00/357) . *Buoy lost. SITE VISIT NOTES: Release was on bottom at depl position, meaning buoy either sank or was vandalized and went adrift.
PM208 - Deployed 01/112 - *. Tube stopped transmitting (01/163). *Buoy lost. SITE VISIT NOTES: Release responded and indicated that it was horizontal.
PM240 - Deployed 01/276- 02/086. All subsurface except SSC failed (01/346). REC NOTES: Raft of wood and netting about 3x5 tied to buoy, cut loose before rec. Lots of guano. Top section was completely disconnected and swinging freely. No bolts, wires torn out, bolt holes did not look overly enlarged.
PM276- Deployed 02/087 - *. Wind vane 0 (02/095). Buoy adrift, then stopped transmitting (02/127). Buoy lost.
PM326 - Deployed 02/290 - 03/304. Rain failed at depl. LWR failed 2003/066.SWAP NOTES: Lots of guano, both rad shields bent, steel ring shackled to bottom of tower leg.
PM326B 2003/108 Swapped LWR and rain. Wind speed and vane 0 2003/136. TC 5 0 2003/108, T20 failed 2003/150, T10 faile 2003/1999. Buoy adrift to NW 2003/199 and out of range for site files 2003/209. REC NOTES: Nylon cut 100-150m above release, definitely by fisherfolk. Anemometer gone, mast had groove melted into it at very top with plastic slag hanging down. SWR shield was bent, SS ring attached to fat leg with remnants of line and on bridle the top section had black line carefully wrapped and tied to the bridle leg just above the weight bucket. Fishing gear at various spots on line. Top 3 modules missing, Sontek was there. Cable looked to have been unscrewed from TCV.

0 95W
PM117 - Deployed 99/330 - 00/310. REC NOTES - Fishing vessel left "thank you" note about their fishing success at this site dated 28 Jan 2000
PM117B - 00/117 Swapped tube, 0-40m modules, added LWR and BP. Wind -> 0 (00/213) , buoy moved ~ 18nm NW. Email from tuna taggers taking CTD's, XBT's near buoy reported mooring had recently been fished (00/259). REC NOTES - 30m long knot in nilspin at 475m. Wind speed bearing frozen, 10m TCV missing.
PM179 - Deployed 00/312 - *. RH failed (00/317).
PM179B - 01/112 - Swapped ATRH. Buoy moved 12 nm and was riding low , estimated scope 0.973. Tube stopped transmitting (01/139) * Buoy lost. No release at deployment or last recorded position.
PM241 - Deployed 01/277 - 01/330. All subsurface 0 (01/292) REC NOTES: Top section broke free from nilspin, wires pulled out of nilspin termination.
PM267 - Deployed 01/330 - 02/290. Used all instrumentation from PM241. SSC failed 02/036
PM267B - 02/085. Added SSC to bridle (divers) w/hose clamps and panduits. Top anti-fouling puck missing from RT SSC, lots of barnacle growth. Unable to use pickle fork to exchange realtime SSC because buoy has insert, new SSC is internally recording but not transmitting. Divers cut fishing line from 12m - 20m. Ring on tower bent, divot out of toroid. Buoy moved 35nm, wind and ATRH failed. REC NOTES: tcv 12 (12432) missing, tc1 11860 missing, tc2 12422 missing. Anemometer missing, stub of wire left behind. Rad mast mount bent ~15 deg on tower ring,
shortwave outer dome smashed. Rad shields bent. Various line attached to bridle and tower. New strobe light on tower. Nilspin jacketing cut through by small line at 7m. Both sontek wires broken, much spiral wrap jammed into a big ball. Longline on top 100m or so.
PM327 - Deployed 02/291-2003/107. Rain failed after depl. V10 failed 2002/312. Wind speed and vane 0 2003/085, LWR failed 2003/067. RH failed 2003/087. Subsurf int 0 at 100m, 120m, 140m, 180m. REC NOTES: 6' loop of hawser on fat leg, ATRH mount bent, anemometer missing nilspin jacketing cut through at 15m, 67m, 100m. Several floats and pieces of longline attached.
PM365 - Deployed 03/107. Buoy moved ~ 6nm, probably fishing boat yant (slingshot). 2003/237 - LWR failed.
PM365B - Removed LWR and SWR. SWAP NOTES: Rad mast was cocked outboard at 35 deg from horizontal, unable to fix so removed sensors. ATRH was bent inboard almost 90 deg. Rain checked out ok.

2S 95W
PM116 - Deployed 99/330 - 00/309. Buoy moved 7nm. Wind -0 (00/048) , RMY broken off below mast, SWR shield bent, rain gauge top broken off above clamp.
PM116B - 00/118 Swapped tube, 0-40m modules, SWR, rain. Added LWR and BP.SSC and SST 0 (00/124) . Wind 0 (00/135) . ATRH failed, Rad/rain failed, (00/195) tube xmitting bad checksums and no good met buffers for 36 days (00/193 -00/229). Email from tuna taggers taking CTD's, XBT's near buoy reported mooring had recently been fished (00/259). REC NOTES -Anemometer missing, rain gauge missing, mast loose from tophat, tower leg broken at weld, rad shield bent, fishing line present from 0 - 100m
PM178 - Deployed 00/310 -01/331. RH too high from dep. Wind failed (00/317).
PM178B - 01/111 - Swapped wind sensor, ATRH, rain gauge. SWAP NOTES: Wind sensor broken off, top of rain gauge broken off. Lines (from fishing vessel) tied to buoy legs. Wind, SWR, LWR failed (01/152). Co/va ok.
PM178C - 01/283 - Swapped wind rain, SWR, LWR. SWAP NOTES: SWR/LWR and platform gone, anemometer gone, mast bent. Rain gauge damaged, replacement not perfectly vertical due to bent bracket REC NOTES: Bent rad mast, bent rain gauge bracket, mast holding to tube with only 1 bolt. TC4, T8, TC11, TCV12 missing.
PM268 - Deployed 01/331 - 02/292.
PM268B - 02/080 -Swapped anemometer. Winds were questionable but nothing obviously wrong with old sensor at recovery. SSC failed and buoy moved to NNW (02/188), appears to be anchored - total move 21nm. REC NOTES: 3" amd 2" line attached to tower. Longline shallow and around nylon spools 3 and 4. SSC cable broken. ATRH bent up 45 deg.LWR radiation shield missing but screws in place. Brain missing, screws gone.
PM328 - Deployed 02/292-03/307. Rains noisy in realtime, RH erratic. V and C 10 intermittantly bad realtime. 2003/048 wind speed and vane 0. SWAP NOTES: . Anemometer had been broken off.
PM328B 2003/106 Swapped wind, rain, atrh. 2003/119, c10 failed. RH maxed and may be bad, was flagged 2003/173. Could be bad entire "B" deployment. C20 consistent density inversion 2003/191. TP 300 appears to have slid down the wire to 500m, flagged 2003/235. All sensors failed 2003/260. Buoy stopped xmitting 2003/281. REC NOTES: No tower or above surface instrumentation. Buoy was 9.2nm north of last position. Tower was broken off at 2 legs with a 2' oiece of 3rd leg remaining. Everything present below buoy except TP300. Hawser was tied to buoy and there was fishing line at various places on the mooring, most notably a wad with a float fouled around the Sontek. Was also some around the 3rd reel near the top

5S 95W
PM84B - Depl 99/328 - 00/119. Rain gauge had top piece above clamp missing and remains were smashed in.
PM144 - Depl 00/119 - 01/110. Wind U/V vane 0. REC NOTES: Wind sensor broken off and missing. Rain gauge top broken off. Modules at 5m, 10m, and 150m missing.
PM206A - Depl 01/110. Wind U/V 0, rain, SWR, RH, BP failed and buoy moved 6.5 nm to NW(01/187).AT low 3 days later. ADRIFT (01/187). Subsurf int 0.REC NOTES: Large wire bridle shackled to tower legs, tower legs worn badly by cable. TC11 (5m) missing. Large hole 4x2ft on underside of toroid. Nylon had parted in 1st reel, about 30m down. Recovered at 3-56.67S, 107-14.99W
PM269 - Deployed 01/332 - 02/297.
PM269B - 02/080 - Swapped anemometer. Old RMY making noise like the prop bearings were worn although the met obs matched the ship well. 02/192 -Buoy adrift. Rec NOTES: TP9 (300m) module missing, blue paint on side of anemometer, 3rd nylon down (E335) parted.
PM329 - Deployed 02/293-*
. Rain noisy in realtime. C10 reads low. *LOST AT SEA. Buoy stopped xmitting hours before the ship arrived for repair. Release located approx 2nm NE of last known buoy position
PM364 - Deployed 03/105- 03/308*-.Buoy stopped xmitting 2003/113 . *LOST AT SEA. Release did not respond

8S 95W
PM115 - Depl 00/119 - 00/306. T/P 500 failed at depl.
PM115B - 00/121 Swapped tube, wind, added LWR and BP. REC NOTES - All ok
PM177A - Deployed 00/307-01/307. SWR and LWR failed (00/330) Subsurface data 0 after Jan 2.
Rain gauge noisy. Stopped xmitting (01/207), no sign of vandalism before failure.
PM177B - Swapped anemometer (no TAO) with hope that tube is recording internally. SWAP NOTES: Anemometer broken, both radiometers missing, rain gauge damaged. Buoy 5nm west of last recorded position. REC NOTES: TC4, T8, T9, TCV 10 and sontek missing. Top 2 shackles were missing cotter pins and nuts. Top ring on tower cracked.
PM252 - Deployed 01/308. T140 failed (01/354). Rain data all 0 02/015
PM252B - 02/079 - Swapped rain gauge, no signs of problems w/old gauge. REC NOTES: All modules and all
surface sensors present and no apparent damage.
PM331 - Deployed 02/295-03/103. Rain noisy in realtime. C10 reads low. REC NOTES: 12m Sontek pulled down ~2m, cable pulled off TCV module.
PM363- Deployed 03/104-03/309 Buoy stopped xmitting 2003/154. REC NOTES: The tower and all its accoutrements were nowhere to be seen. Only the stub of one leg left bolted to the buoy. The 2 others were missing bolts. The underside of the buoy was abraded and cracked, probably by a wire rope tow line or tie up line. The SSC was gone, bolt and all. Only 2 small pieces of the module mount left where it was screwed to the SSC plate. All other modules present, but the 60 meter module was not inverted (it had 2 module mounts and was noted on the deployment log as having been inverted). Two other modules (140, 180) were completely out of their mounts 180 degrees, one being held on by one panduit and in both cases the module mount was broken. Other than that recovery was normal.

TAO/EPIC Cruises along 95W

Dates Ship Ports Cruise CS
24 Oct - 25 Nov Ron Brown Panama - Panama GP6-03-RB Ben Moore
24 Mar - 28 Apr Ka'imimoana San Diego - Puntarenas GP2-03-KA Patrick A'Hearn
8 Oct- 8 Nov 02 Ron Brown Panama - Panama GP6-02-RB Patrick A'Hearn
1 Mar - 4 Apr 02 Ka'imimoana San Diego - Manzanillo GP1-02-KA Mike McPhaden / Ben Moore
29 Oct - 05 Dec 01 Ron Brown Chile- Panama GP8-01-RB Brian Lake
05 Sep - 06 Oct 01 Ron Brown San Diego- Galapagos GP6-01-RB *John Shanley
28 Mar - 05 May 01 Ka'imimoana San Diego - San Diego GP2-01-KA Dave Zimmerman
16 Oct - 16 Nov 00 Ron Brown San Diego - Panama GP7-00-RB Rick Miller
12 Apr - 21 May 00 Ka'imimoana San Diego - San Diego GP2-00-KA Ben Moore
6 Nov - 9 Dec 99 Ron Brown San Diego - San Diego GP8-99-RB Andy Shepherd
*TAO lead, not chief scientist
TAO EPIC Home | Research | QuickLook95W | Data display | Data delivery | TAO Home

TAO Project Office
| Pacific Marine Environmental Laboratory
7600 Sand Point Way NE
Seattle, WA 98115
Disclaimer | Privacy Policy